one of the angles of a triangle has a measure of 40degrees. find the measures of the other two angles if the difference between their measures is 10degrees
Answer by ewatrrr(24785)

180-40
140 degrees left for the other two
x + (x-10) = 140
2x = 150
x = 75 and (x-10) = 65
And...
40 + 75 + 65 = 180